OpenStudy (calculusxy):

The position of a bug is described by the parametric equation (x, y) = (2 − 12t, 1 + 5t). Explain why the speed of the bug is 13 cm/sec. Change the equation to obtain the description of a bug moving along the same line with speed 26 cm/second.

OpenStudy (calculusxy):

@zepdrix @agent0smith @skullpatrol

OpenStudy (agent0smith):

Notice... x = 2 - 12t <= in one second, it advances 12 units in the negative x direction y = 1 + 5t <= in one second, it advances 5 units in the positive y direction 

OpenStudy (calculusxy):

OH. It creates 5-12-13 triangle!

OpenStudy (calculusxy):

Now to have the speed be 26cm/sec do I need to double 5 and 12?

OpenStudy (agent0smith):

Yes. Changing those shouldn't affect the line at all... I don't think. You might have to double everything in both equations. We can check: x = 2 - 12t <==== solve for t y = 1 + 5t <==== plug in here x = 4 - 24t <==== solve for t y = 2 + 10t <==== plug in here Do you get the same equation? If not, then it's not moving along the same line. If that fails, just try doubling only the 12 and 5 and see if that yields the same line as the original.

OpenStudy (calculusxy):

OK I am trying it out right now

OpenStudy (calculusxy):

Well the slope is basically the same (if you completely simplify it). But do you think that I should make a right triangle and solve for the hypotenuse and see if it is 26?

OpenStudy (calculusxy):

Just checked for the hypotenuse and it is 26.

OpenStudy (calculusxy):

@agent0smith Can you help me with another question please?

OpenStudy (agent0smith):

But you need to check it's the exact same line...

OpenStudy (calculusxy):

So you are saying to solve for t and then plug it back in to the parametric equation for the y. How would I solve for t? (This might be a dumb question)

OpenStudy (agent0smith):

You should know how to solve for t. Get it alone by doing math.

OpenStudy (calculusxy):

I am so sorry but I am really bad with parametric equations. But if I am wrong with this please correct me. So would the x value be like 0 in the coordinate (0,5)

OpenStudy (agent0smith):

x = 2 - 12t rearrange to solve for t

OpenStudy (calculusxy):

Or... 2-12(1) = 2 - 12 = -10 <- x 1 + 5(1) = 1 + 5 = 6 <- y

OpenStudy (calculusxy):

No that doesn't make any sense.

OpenStudy (calculusxy):

What do you mean by "rearrange to solve for t"?

OpenStudy (agent0smith):

Get t alone. so t = something.

OpenStudy (calculusxy):

2-12t = x -12t = x - 2 t = (x/(-12)) + 1/6

OpenStudy (calculusxy):

Is that right?

OpenStudy (agent0smith):

Yes. Now you plug that into the y.

OpenStudy (calculusxy):

y = 1 + 5((x/(-12)) + 1/6) y = 1 + (5x/(-12) + 5/6) y = (1 5/6) + (5x/(-12))

OpenStudy (agent0smith):

Now you gotta make sure the second gives the same line.

OpenStudy (calculusxy):

ok

OpenStudy (calculusxy):

4 - 24t = x -24t = x - 4 t = (x/(-24) + 1/6) y = 2 + 10(x/(-24) + 1/6) y = (2 1/6) + (5x/(-12))

OpenStudy (calculusxy):

2 - 24t = x 1 + 10t = y -24t = x - 2 t = (x/(-24) + 1/12) 1 + 10((x/(-24) + 1/12)) 1 + 10x/(-24) + 10/12 1 + 5x/(-12) + 5/6 1 5/6 + (5x - 12) <- this works!

OpenStudy (agent0smith):

\[\checkmark \]

OpenStudy (calculusxy):

Can I ask you like two more questions please?

OpenStudy (calculusxy):

AND THANK YOU!

OpenStudy (agent0smith):

I gotta go, i'll be back later.

OpenStudy (calculusxy):

Ok thank you for your help though :)